40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 13 17:37:43 2014, Jenne, Update, General, Game plan 2014_Aug_13.pdf
    Reply  Thu Aug 14 02:51:46 2014, Jenne, Update, General, Game plan Y_ALS_FINE_PHASE_OUT_noisy_13Aug2014.pdf
    Reply  Thu Aug 14 14:58:03 2014, Jenne, Update, General, Updated game plan 2014_Aug_14.pdf
       Reply  Thu Aug 14 15:42:29 2014, Koji, Update, General, Updated game plan 
          Reply  Thu Aug 14 15:51:37 2014, Jenne, Update, General, Updated game plan 
             Reply  Thu Aug 14 18:02:11 2014, Koji, Update, General, Updated game plan 
                Reply  Thu Aug 14 18:05:05 2014, Jenne, Update, General, Updated game plan 
    Reply  Fri Aug 15 20:26:37 2014, ericq, Update, General, Game plan 
Message ID: 10382     Entry time: Thu Aug 14 02:51:46 2014     In reply to: 10377
Author: Jenne 
Type: Update 
Category: General 
Subject: Game plan 

[Jenne, Rana]

* Decided that earlier mode hop scan won't give us the information that we were hoping for.  We need to think about where we can actually see the frequency change.  Can we use the IR beatnote that we will soon have to do this?  We'd only be able to scan one laser temp at a time, but that's okay.  Leave, say, the PSL temperature alone, and scan one of the end laser temps.  Using the PSL as the reference, we will be able to see if the frequency of the end laser goes crazy and jumpy as we pass through a certain temp.  Then, repeat while holding the end laser constant and scan the PSL.  Thoughts?

* Meditated on PSL oplev servo, but I need to make a Matlab script that can evaluate different loops according to a cost function based on elog 9690

* Aligned IFO to IR, then greens to arms (got back to 0.9 for GTRY, but only about 0.5 for GTRX, with the PSL green shutter closed).  Then aligned green beams on the PSL table, since the PSL green pointing had changed a bit from Q's crystal alignment tweak-up earlier today.  Beatnotes are nice and big (see elog 10381 - The Yarm is the larger beatnote, and the Xarm is the smaller one.)

* Was not able to lock ALS comm/diff and hold long enough to get both arms to IR resonance.  Also, saw that TRY's RIN was more than 50%(!!!).  We took a look, and there seems to be much more low frequency noise than there was when the spectrum in the control room was taken for the multicolor metrology paper:

Y_ALS_FINE_PHASE_OUT_noisy_13Aug2014.pdf

* Tried to balance the ALS comm/diff input matrix, with not a lot of success.  First of all, it looks like the Xarm has overall about 10 times more noise!  We were exciting MC2 in position (~88 Hz, about 130 counts I think), and then looking at DARM_IN1 for the peak.  When DARM_IN1 was just one of the 2 ALS error signals (i.e. one matrix element set to zero), versus when both matrix elements were set to 1, we saw a factor of only about 3 in reduction of the peak height.  We were hoping to have better cancellation of this pure CARM signal in the DARM channel.  The Xarm green PDH loses lock every ~5 or 10 minutes, and when we relock it, this cancellation seems different, so we want to try again tomorrow when the ALS is locked on comm / diff, rather than just the free running ALS that we have now.  Although, if the balance of the input matrix changes lock-to-lock, we may need to consider redoing the green PSL table layout so we get a pure DARM beatnote signal like they have at the sites.

* We want to change how the watch script for ALS works, although this is a low-priority task.  Rather than looking at the control signal, we should maybe look at the sum of all the coil outputs, multiplied by a pendulum TF, and use that as a rough displacement sensor.  We want to be careful of pushing too hard at low frequencies, but we want to allow higher frequency actuation without having the watch script shut things down.

* Also, I should put on the to-do list the revamp of the ALS find IR resonance script.

ELOG V3.1.3-